OpenStudy (anonymous):

Looking for the correct following problem where 3 of the 4 knowns provided. I have a liquid sample that weighs 8.9#/gallon. (the product is a fermented product) If I start with 100 gallons of this product and end up with 8# of dry product..what is the amount of dry product if I start with 2.5 gallons(weighing the same)?

OpenStudy (anonymous):

Simple proportionnality, methink : \[\frac{a}{b} = \frac{c}{x} \Leftrightarrow x = \frac{b \times c}{a}\] So : \[\frac{100}{8} = \frac{2.5}{x} \Leftrightarrow x = \frac{8 \times 2.5}{100} \Leftrightarrow x = 0.2 \] The starting quantity is *0.2#